A car has an acceleration of 2.4m/s assume that its acceleration remains constant, starting from rest how long does the car require to reach a velocity of 90 km/h ? How far does it travel while reaching that velocity ?




1
Expert's answer
2021-11-17T10:37:26-0500

(a)

"v=v_0+at,""t=\dfrac{v-v_0}{a}=\dfrac{90\ \dfrac{km}{h}\times\dfrac{1000\ m}{1\ km}\times\dfrac{1\ h}{3600\ s}-0}{2.4\ \dfrac{m}{s^2}}=10.42\ s."

(b)

"v^2=v_0^2+2ad,""d=\dfrac{v^2}{2a}=\dfrac{(90\ \dfrac{km}{h}\times\dfrac{1000\ m}{1\ km}\times\dfrac{1\ h}{3600\ s})^2}{2\times2.4\ \dfrac{m}{s^2}}=130.21\ m."